# Active Light Modulation to Counter Manipulation of Speech Visual Content

### v1

@@ A.4. Perceptibility Evaluation Details
User study   We invite groups of 9 participants (recruited via email, with ages ranging from 18 to 34) to each of the test environments (Figure 9). At each environment participants were asked to assess the projection surface during four 45 s trials in which the core unit either performed embedding of a random bitstream or was powered off as a control scenario (done for two randomly selected trials). Participants were informed that light may be projected and shown the projection region boundaries. During each trial they were allowed to walk freely up to 1.5 m of the surface, and asked to answer two questions once ready:
@@
As shown in the bottom two figures of Figure 11, for each environment we plot (1) $\Delta=TPR-FPR$, where TPR is the rate at which participants responded ”Yes” to Q1 when embedding was occurring, and FPR is the rate at which they incorrectly responded ”Yes”; and (2) the average response to Q2, excluding those given with a Q1 false positive. A $\Delta$ value $\leq 0$ indicates that embedding is effectively imperceptible, as participants perform no better than random at detecting it. We observe this in all but two video cases. In person, $\Delta\leq 0.2$ in all but two environments. Respondents uniformly report low obtrusiveness in-video and live. The primary factors influencing perceptibility are the surface’s texture and color. Darker, homogeneous surfaces (i.e., S2, S4, S5) fundamentally contrast with impinging light, whereas brighter-colored, more complex ones (e.g, S1, S3, S5) provide a camouflage. Increasing ambient light intensity can counteract this (as with S4 and S2 at 750 lx) by increasing the baseline brightness of a surface’s appearance.
Perceptual metrics   We additionally evaluate optical signature perceptibility in video using the learned perceptual loss (LPIPS) [149, 154] metric. This metric has been shown to have superior correlation with human perception, especially in the context of fine-grained changes such as those introduced by Spotlight’s optical modulations. LPIPS takes as input two images and outputs a score ranging from 0 to 1, where a lower value indicates the inputs are more perceptually similar. In particular, prior works show humans cannot sense differences when LPIPS is below 0.5 [155]. To apply these image-level metrics to our videos, we compute the average score between crops of the full projection region and individual cells in 5,000 pairs of frames captured with and without Spotlight operating. We compute at both the full projection region and cell level to understand perceptibility at both scales differences occur.
@@
Figure 11. Embedding perceptibility results. Top axis: Average LPIPS across pairs of video frames captured w/ and w/o Spotlight operation. All scores are significantly below the established LPIPS perceptibility threshold of 0.5 [155]. Middle: average obtrusiveness reported by users study respondents. Bottom: user study $\Delta=TPR-FPR$, normalizing true positive identifications of embedding occurrence to false positive identifications during control trials. A nonpositive value indicates imperceptibility.
A.5. Proofs and Definitions: Locality-Sensitive Hashing
@@
Here we detail our methodology for utilizing cosine similarity-based LSH for verification. We include both a definition and discussion of the methodology at a high level (Definition 2) and derive an equation for the relationship between hash size and verification performance (Theorem 4), which illustrates that verification performance has no dependence on input vector dimensionality.
@@
We utilize the locality sensitive hashing scheme for the cosine similarity as proposed in [88]. Let $\vec{r}$ be a vector in $\mathbb{R}^{n}$ chosen randomly from the n-dimensional Gaussian distribution (i.e., each coordinate is drawn from a Gaussian distribution). Let the hash function $h_{\vec{r}}:\vec{u}\in\mathbb{R}^{n}\mapsto\{0,1\}$ be defined as follows:
@@
The locality sensitive hashing scheme $H$ is defined as $H(\vec{u})=\{h_{1}(\vec{u}),h_{2}(\vec{u}),...,h_{k}(\vec{u})\}$, where $h_{i}$ are independently and randomly chosen hash functions from $\mathbb{F}$. Thus, given an input $\vec{u}\in\mathbb{R}^{n}$, $H$ outputs a $k$-bit vector, formed by concatenating the single bit outputs of each of its hash functions $h_{i}$.
The key idea of this scheme is that the sign of a vector $\vec{x}$’s projection onto $\vec{r}_{i}$ is fundamentally related to the angle between $vec{x}$ and $\vec{r}_{i}$. Thus if $\vec{u}$ and $\vec{v}$ have a high cosine similarity, their projections onto $\vec{r}_{i}$ are more likely to have the same sign. Each bit in the hash serves as an additional “sample” to aid in approximating $\Theta$; including more bits increases the probability that $D$ correctly reflects $\Theta$. This captures the cosine similarity with extreme space efficiency.
@@
We depart from the formal definition of a traditional verification problem: given two feature vectors, we would like to confirm that they represent the same source (e.g., face embeddings corresponding to the same identity). Two feature vectors $u,v$ are said to correspond to the source if $\Theta(\vec{u},\vec{v})<\theta_{th}$, where $\theta_{th}$ is a decision threshold. Otherwise, the vectors are said to correspond to different sources.
We can similarly formalize the verification problem on hashed feature vectors: the hashes of two feature vectors, $H_{cos}(u),H_{cos}(v)$ are said to correspond to the same source if $D(H(\vec{v}),H(\vec{u}))>d_{th}$, where $d_{th}$ is a decision boundary for verification on hashed vectors. If $\theta_{th}$ is the optimal decision boundary for verification on the raw feature vectors, intuitively, the optimal value of $d_{th}$ should be the expected value of $D(H(\vec{v}),H(\vec{u}))$ for two vectors $u,v$ s.t. $\Theta(\vec{u},\vec{v})=\theta_{th}$. From Equation 3, this is $\frac{k\theta_{th}}{\pi}$.
@@
Recall that $D$, the Hamming distance function, gives the number of positions at which the values of two bitstrings differ. Since the value of each bit of $H$’s output is determined by a hash function $h_{i}$, the expected value of $D(H(\vec{v}),H(\vec{u}))$ is the expected number of $H$’s $k$ hash functions $h_{i}$ for which $h_{i}(\vec{u})\neq h_{i}(\vec{v})$. From [88],
@@
(2) (${k\choose n}\mathbb{Pr}[h(\vec{u})\neq h(\vec{v})]^{n}\mathbb{Pr}[h(\vec{u})=h(\vec{v})]^{n-k}$
=($={k\choose n}\left(\frac{\Theta(\vec{u},\vec{v})}{\pi}\right)^{n}(1-\frac{\Theta(\vec{u},\vec{v})}{\pi})^{k-n}$
@@
To assess the impact of the hash size $k$ on verification performance, we seek the probability $P_{\theta_{th}}(k)$ that all decisions obtained from verification on $k$-bit hashed feature vectors are the same as those obtained from performing verification on the raw vectors with a decision boundary of $\theta_{th}$. This event indicates that the hashed feature vectors perfectly preserve verification performance. Thus we can view the probability of its occurrence as a measure of the hashed vectors’ performance relative to that of the raw vectors. $P_{\theta_{th}}(k)$ is given by the following equation
@@
[…]
@@
Based on Definition 2, verification decisions are obtained from raw or hashed feature vectors using the criteria $\Theta(\vec{u},\vec{v})<\theta_{th}$ or $D(H(\vec{v}),H(\vec{u}))<\frac{k\theta_{th}}{\pi}$, respectively. Thus we have
@@
$\forall(\vec{u},\vec{v})\in\{(\vec{u},\vec{v}):\Theta(\vec{u},\vec{v})>\theta_{th}\}\hskip 8.5359ptD(H(\vec{v}),H(\vec{u}))>\frac{k\theta_{th}}{\pi}$
@@
[…]
@@
[…]
@@
[…]

### v2

@@ A.1. Perceptibility Evaluation Details
User study   We invite groups of 9 participants (recruited via email, with ages ranging from 18 to 34) to each of the test environments (Figure 9). At each environment participants were asked to assess the projection surface during four 45 s trials in which the core unit either performed embedding of a random bitstream or was powered off as a control scenario (done for two randomly selected trials). Participants were informed that light may be projected and shown the projection region boundaries. During each trial they were allowed to walk freely up to 1.5 m of the surface. They were asked to answer two questions once ready:
@@
As shown in the bottom two figures of Figure 11, for each environment we plot (1) $\Delta=TPR-FPR$, where TPR is the rate at which participants responded ”Yes” to Q1 when embedding was occurring, and FPR is the rate at which they incorrectly responded ”Yes”; and (2) the average response to Q2, excluding those given with a Q1 false positive. A $\Delta$ value $\leq 0$ indicates that embedding is effectively imperceptible, as participants perform no better than random at detecting it. We observe this in all but two video cases. In person, $\Delta\leq 0.2$ in all but two environments. Respondents uniformly report low obtrusiveness in video and live. The primary factors influencing perceptibility are the surface’s texture and color. Darker, homogeneous surfaces (i.e., S2, S4, S5) fundamentally contrast with impinging light, whereas brighter-colored, more complex ones (e.g, S1, S3, S5) provide a camouflage. Increasing ambient light intensity can counteract this (as with S4 and S2 at 750 lx) by increasing the baseline brightness of a surface’s appearance.
Perceptual metrics   We additionally evaluate optical signature perceptibility in video using the learned perceptual loss (LPIPS) (Zhang et al., 2018a; pyt, 2025) metric. This metric has been shown to have superior correlation with human perception, especially in the context of fine-grained changes such as those introduced by VeriLight’s optical modulations. LPIPS takes as input two images and outputs a score ranging from 0 to 1, where a lower value indicates the inputs are more perceptually similar. In particular, prior works show humans cannot sense differences when LPIPS is below 0.5 (Laidlaw et al., 2020). To apply these image-level metrics to our videos, we compute the mean LPIPS between crops of the full projection region and individual cells in 5,000 pairs of frames captured with and without VeriLight operating. We compute scores at both the full projection region and cell level to understand perceptibility at both scales viewers may notice differences.
@@
Figure 11. Embedding perceptibility results. Top: Mean LPIPS for pairs of video frames captured w/ and w/o VeriLight operation. All scores are significantly below the established LPIPS perceptibility threshold of 0.5 (Laidlaw et al., 2020). Middle: mean obtrusiveness reported by users study participants. Bottom: user study $\Delta=TPR-FPR$, normalizing true positive identifications of embedding occurrence to false positive identifications during control trials. A nonpositive value indicates imperceptibility.
[…]
% Window
Modified
Meso4
(Afchar et al., 2018)
Xception
(Rossler et al., 2019)
Capsule
(Nguyen et al., 2019)
Efficient
(Tan and Le, 2019)
aSRM
(Luo et al., 2021)
aSPSL
(Liu et al., 2021)
aRecce
(Cao et al., 2022)
aUCF
(Yan et al., 2023a)
aTALL
(Xu et al., 2023)
AltFreeze.
(Wang et al., 2023a)
Ours
10-20 0.53 0.59 0.57 0.57 0.55 0.55 0.55 0.60 0.55 0.47 0.72
20-30 0.58 0.62 0.55 0.58 0.55 0.58 0.55 0.62 0.57 0.47 0.84
30-40 0.60 0.64 0.58 0.60 0.58 0.60 0.56 0.63 0.58 0.47 0.90
40-50 0.59 0.67 0.59 0.63 0.58 0.61 0.59 0.66 0.61 0.48 0.92
100 0.75 0.90 0.80 0.81 0.76 0.76 0.83 0.84 0.84 0.26 0.98
[…]
Table 3. Comparison of AUC scores achieved by passive detectors and VeriLight’s 150-bit dynamic feature hashes on our multi-posed, fine-grained reenactment dataset (§9.1), in which modifications of various granularities (by percentage duration modified) are applied to each 4.5 s window. Best performing method is bolded.
A.2. Adversarial Deepfake Model Training
DaGAN   The DaGAN reenactment model forward pass takes as input two face images: a source image of the victim and an attacker-provided ”driving” image. It synthesizes a fake image where the victim possesses a new facial expression and pose, matching those in the driving image. A video can be formed by performing a forward pass of the model for each pair of frames in an original video and driving video.
During training of our modified DaGAN model, we extract the 16 FaceMesh features considered by VeriLight’s dynamic feature vectors (§4.2) from each source and generated fake image. Our loss function $L_{adv}$ is then defined as follows:
$L_{adv}=L+\alpha\Theta(Dyn_{src},Dyn_{fake})$
where $L$ is the original DaGAN loss function, $Dyn\in\mathbb{R}^{16}$ are vectors containing the FaceMesh results, and $\Theta$ is the cosine similarity function. The coefficient $\alpha$ weights the importance of our adversarial spoofing objective.
To implement this modified version of DaGAN, we first develop a fully differentiable PyTorch implementation of FaceMesh. [^7] We release our MediaPipe FaceMesh implementation with our other code, as linked on our project website This is needed because FaceMech is only distributed as a LiteRT module, which is only designed for inference and does not have the necessary mechanisms for backpropagation. We use ONNX-converted (onn, 2025) ports of each of the three neural networks underlying FaceMesh and integrate them to the best of our abilities by referencing MediaPipe’s public model cards (med, 2024). Our implementation outputs facial landmarks with an average difference of 2 pixels and blendshape scores with an average difference of 0.11 (arbitrary units, ranging from 0 to 1) from their official implementation counterparts when run on DaGAN’s training dataset.
We train our adversarial DaGAN model by fine-tuning with our loss component added. We start training from the checkpoint released by the authors (dag, 2023), applying early stopping based on validation loss. We use the same learning rate and parameters employed in their original implementation, as well as their same training data, sourced from VoxCeleb. We create evaluation videos by randomly choosing 55 pairs of videos from the VoxCeleb test split and using one to drive the other.
FSGAN   The FSGAN identity swap model takes as input a source image of the victim and an attacker-provided target face image. The model synthesizes an image in which the victim’s face is supplanted with the target’s face, effectively modifying the portrayed identity.
During training of our modified FSGAN models, we extract the ArcFace embedding from both the source and generated fake images. Similar to the case of DaGAN, our loss function $L_{adv}$ is then defined as follows:
$L_{adv}=L+\alpha\Theta(Arc_{src},Arc_{fake})$
where $L$ is the original FSGAN loss function, $Arc\in\mathbb{R}^{512}$ are the extracted ArcFace embeddings, and $\Theta$ is the cosine similarity function. The coefficient $\alpha$ again weights the importance of our adversarial spoofing objective.
We train the FSGAN model from scratch using data from the VoxCeleb dataset, because the authors do not release all weights necessary for fine-tuning. We use the same learning parameters employed in the original version and apply early stopping based on validation loss. For testing, we randomly choose 55 pairs of videos from the VoxCeleb dataset (ensuring they portray different identities) and generate an identity swap deepfake for each.
A.3. Proofs and Definitions: Locality-Sensitive Hashing
@@
Here we formalize our approach to using cosine similarity-based LSH for verification. We include both definitions and discussions of the methodology (Definition 2) and derive an equation for the relationship between hash size and verification performance (Theorem 4). This equation illustrates that verification performance has no dependence on input vector dimensionality.
@@
Recall from §4.1 that the LSH scheme for cosine similarity, $H_{cos}$, outputs $k$-bit hashes such that $D(H_{cos}(\vec{u}),H_{cos}(\vec{v})$ estimates $\Theta(\vec{u},\vec{v})$. $\Theta(\vec{u},\vec{v})$ is the angle between $\vec{u}$ and $\vec{v}$, and $D$ is the Hamming distance. As defined in (Charikar, 2002), $H_{cos}$ is composed of multiple random projection hash functions $h$, each outputting one bit.
Specifically, let $\vec{r}$ be a vector in $\mathbb{R}^{n}$ chosen randomly from the $n$-dimensional Gaussian distribution (i.e., each coordinate is drawn from a Gaussian distribution). Let the hash function $h_{\vec{r}}:\vec{u}\in\mathbb{R}^{n}\mapsto\{0,1\}$ be defined as follows:
@@
The locality sensitive hashing scheme $H$ is defined as $H(\vec{u})=\{h_{1}(\vec{u}),h_{2}(\vec{u}),...,h_{k}(\vec{u})\}$, where $h_{i}$ are independently and randomly chosen hash functions of the form above. Thus, given an input $\vec{u}\in\mathbb{R}^{n}$, $H$ outputs a $k$-bit vector, formed by concatenating the single bit outputs of each of its hash functions $h_{i}$.
The key idea of this scheme is that the sign of a vector $\vec{u}$’s projection onto $\vec{r}_{i}$ is fundamentally related to the angle between $\vec{u}$ and $\vec{r}_{i}$. Thus if $\vec{u}$ and another vector $\vec{v}$ have a high cosine similarity, their projections onto $\vec{r}_{i}$ are more likely to have the same sign. Each bit in the hash is an additional instance of this process and “sample” to aid in approximating $\Theta$; including more bits increases the probability that $D$ correctly reflects $\Theta$. This captures the cosine similarity with extreme space efficiency.
@@
We depart from the definition of a traditional verification problem: given two feature vectors, we would like to confirm that they represent the same source (e.g., face embeddings corresponding to the same identity). Two feature vectors $u,v$ are said to correspond to the source if $\Theta(\vec{u},\vec{v})\leq\theta_{th}$, where $\theta_{th}$ is a decision threshold. Otherwise, the vectors are said to correspond to different sources.
We can similarly formalize the verification problem on hashed feature vectors: the hashes of two feature vectors, $H_{cos}(u),H_{cos}(v)$ are said to correspond to the same source if $D(H(\vec{v}),H(\vec{u}))\leq d_{th}$, where $d_{th}$ is a decision threshold for verification on hashed vectors. If $\theta_{th}$ is the optimal decision threshold for verification on the raw feature vectors, intuitively, the optimal value of $d_{th}$ should be the expected value of $D(H(\vec{v}),H(\vec{u}))$ for two vectors $u,v$ s.t. $\Theta(\vec{u},\vec{v})=\theta_{th}$. From Theorem 3 below, this is $\frac{k\theta_{th}}{\pi}$.
@@
Recall that $D$, the Hamming distance function, gives the number of positions at which the values of two bitstrings differ. Since the value of each bit of $H$’s output is determined by a hash function $h_{i}$, the expected value of $D(H(\vec{v}),H(\vec{u}))$ is the expected number of $H$’s $k$ hash functions $h_{i}$ for which $h_{i}(\vec{u})\neq h_{i}(\vec{v})$. From (Charikar, 2002),
@@
(2) ${k\choose n}\mathbb{Pr}[h(\vec{u})\neq h(\vec{v})]^{n}\mathbb{Pr}[h(\vec{u})=h(\vec{v})]^{n-k}$
$={k\choose n}\left(\frac{\Theta(\vec{u},\vec{v})}{\pi}\right)^{n}(1-\frac{\Theta(\vec{u},\vec{v})}{\pi})^{k-n}$
@@
To assess the impact of the hash size $k$ on verification performance, we seek the probability $P_{\theta_{th}}(k)$ that all decisions obtained from verification on $k$-bit hashed feature vectors are the same as those obtained from performing verification on the raw vectors with a decision threshold of $\theta_{th}$. This event indicates that the hashed feature vectors perfectly preserve verification performance. Thus we can view the probability of its occurrence as a measure of the hashed vectors’ performance relative to that of the raw vectors. $P_{\theta_{th}}(k)$ is given by the equation
@@
$\exp\left(\int_{\theta_{th}}^{\pi}\ln\left(\sum_{n=\frac{k\theta_{th}}{\pi}}^{k}{k\choose n}\left(\frac{\theta}{\pi}\right)^{n}\left(1-\frac{\theta}{\pi}\right)^{k-n}\right)d\theta\right)$
Figure 12. $\bm{P_{\theta_{th}}(k)}$ (derived in Theorem 2) plotted for various cosine similarity decision thresholds $\bm{\theta_{th}}$. Our results are obtained using a cosine similarity threshold of 1.17 for our raw dynamic feature vectors and 0.88 for our ArcFace identity feature vectors (§4.2). Notably, all curves’ elbows occur before our chosen hash size $\bm{k=150}$.
Evidently, none of the equation’s terms depend on the dimensionality of the input vectors. $P_{\theta_{th}}(k)$ is only a function of the decision threshold $\theta_{th}$ and $k$.
@@
Based on Definition 2, verification decisions are obtained from raw or hashed feature vectors using the criteria $\Theta(\vec{u},\vec{v})\leq\theta_{th}$ or $D(H(\vec{v}),H(\vec{u}))\leq\frac{k\theta_{th}}{\pi}$, respectively. Thus we have
@@
$\hskip-28.45274pt\forall(\vec{u},\vec{v})\in\{(\vec{u},\vec{v}):\Theta(\vec{u},\vec{v})>\theta_{th}\}\hskip 8.5359ptD(H(\vec{v}),H(\vec{u}))>\frac{k\theta_{th}}{\pi}$
@@
$=\exp\left(\int_{0}^{\theta_{th}}\ln\left(\sum_{n=0}^{\frac{k\theta_{th}}{\pi}}{k\choose n}\left(\frac{\theta}{\pi}\right)^{n}\left(1-\frac{\theta}{\pi}\right)^{k-n}\right)d\theta\right)$
@@
$=\exp\left(\int_{\theta_{th}}^{\pi}\ln\left(\sum_{n=\frac{k\theta_{th}}{\pi}}^{k}{k\choose n}\left(\frac{\theta}{\pi}\right)^{n}\left(1-\frac{\theta}{\pi}\right)^{k-n}\right)d\theta\right)$
@@
$\exp\left(\int_{\theta_{th}}^{\pi}\ln\left(\sum_{n=\frac{k\theta_{th}}{\pi}}^{k}{k\choose n}\left(\frac{\theta}{\pi}\right)^{n}\left(1-\frac{\theta}{\pi}\right)^{k-n}\right)d\theta\right)$